A bat can detect small objects, such as an insect, whose size is approximately equal to one wavelength Of sound the bat makes, if bat's emit a chirp at a frequency of 6.00 × 10⁴ Hz, and if the speed of sound in air is 343m/s, what is the smallest insect a bat can detect?
1
Expert's answer
2020-06-08T10:25:23-0400

If the frequency is "f" and the speed of sound is "v" , then the wavelength can be calculated as (see https://en.wikipedia.org/wiki/Wavelength)

"\lambda = \dfrac{v}{f} = \dfrac{343\,\mathrm{m/s}}{6.00\times10^4\,\mathrm{Hz}} = 5.7\cdot10^{-3}\,\mathrm{m} = 5.7\,\mathrm{mm}."